OpenStudy (anonymous):

What is the value of the expression |x| + |y + z| when x = –7, y = 8, and z = –11?

OpenStudy (anonymous):

plug in the values and evaluate

OpenStudy (anonymous):

\(|-7|+|8+(-11)|\)

OpenStudy (anonymous):

what's the absolute value of -7?

OpenStudy (anonymous):

what's 8 + (-11) = ? what's the absolute value of that result?

OpenStudy (anonymous):

19?

OpenStudy (anonymous):

nope... 8 + (-11) = 8 + [(-8) + (-3)] = [8 + (-8)] + (-3) = 0 + (-3) = -3 what's the absolute value of -3?